Required information The following information applies to the questions displayed below. ELS Corporation is about to begin its sixth year of existence. Assume that ELS reported gross receipts for each of its first five years of existence for Scenarios A, B, and C as follows: Year of Existence Scenario A Scenario B Scenario 1$ 20,000,000 $15,000,000 $24,500,000 2 25,000,000 25,000,000 25,000,000 60,000,000 35,000,000 30,000,000 80,000,000 35,000,000 40,000,000 100,000,000 40,000,000 55, 000,000 c. In what years is ELS allowed to use the cash method of accounting under Scenario C? Year 1, 2, and 3 Year 1, 3, 4, and 5 Year 1, 2, 3, and 4 Year 3,4, and 5 Year 1, 5, and 6
Answer #1
Cash method of accounting can be used in years in which average gross receipts for the past three years is $25 million or less
Year Average Gross Receipts
1 -
2 24500000
3 24750000 =(24500000+25000000)/2
4 26500000 =(24500000+25000000+30000000)/3
5 31666667 =(25000000+30000000+40000000)/3
6 41666667 =(30000000+40000000+55000000)/3
ELS can use cash method of Accounting in Year 1,2 and 3
Option 1 Year 1,2,and 3 is correct